OpenStudy (anonymous):

determine whether the lines x = 3t+2, y= t-1, z = 6t+1, and x= 3s-1. y= s-2, z = s intersect.

OpenStudy (anonymous):

First you factor in the quantum mechanics. Here's my work.

OpenStudy (anonymous):

(7,9)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

HAhaha IStutts is a NIMROD!!!

OpenStudy (anonymous):

i do not know quantum mechanics

OpenStudy (anonymous):

You don't get it?

OpenStudy (anonymous):

reminds me a little of a matrix, but that's it

OpenStudy (anonymous):

Well, then you shouldn't be taking a quantum mechanics class?

OpenStudy (anonymous):

didn't know vector calc was quontum mechanics

OpenStudy (anonymous):

quantum*

OpenStudy (anonymous):

Sooooo, I guess that's the only way you know to solve that?

OpenStudy (anonymous):

Yeah, some reason in your calc class, quantum mechanics is involved.

OpenStudy (anonymous):

well, i appreciate the help. haha! at least you didnt call me a nimrod

OpenStudy (anonymous):

Any takers?

OpenStudy (anonymous):

ok solve the following set the x and components equal 3t+2=3s-1 t-1=s-2 solve for s and t let me know what you get.

OpenStudy (anonymous):

@IStutts can you solve these two equations ? 3t+2=3s-1 t-1=s-2